Dylan drove to the beach,which is one hundred thirty-six kilometers away,at forty-eight kph. He drove home the next day at sixty kph. How many hours in total did Dylan spend driving?
Answer by checkley77(12569)

136/48+136/60=2.833+2.267=5.1 hours for the round trip.